OpenStudy (anonymous):

Jenny Is solving the equation x^2-8x=9 by completing the square. What number should be added to both sides of the equation to complete the square?

OpenStudy (anonymous):

$$(\frac b 2)^2$$

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

b=8 1/2b=4 (1/2b)^2=16

OpenStudy (anonymous):

ok, thank-you

OpenStudy (anonymous):

I though that we had to plug in a number for one of the variables

OpenStudy (anonymous):

but, I get it now, thank-you